OpenStudy (anonymous):

What is the value of the y variable in the solution to the following system of equations? 3x + y = 11 -3x + 2y = 1 2 4 y can be any number as there are infinitely many solutions to the system. There is no y value as there is no solution to the system.

OpenStudy (lasttccasey):

Well the easiest way to solve this would be to solve the first equation for x where y is the variable and plug that into the second equations so you will instantly solve for y.

OpenStudy (anonymous):

Can you walk me through it

OpenStudy (anonymous):

@lasttccasey

OpenStudy (lasttccasey):

Sure, first take the first equation, 3x + y = 11, and solve it for x. So make this equation in the form x = ...

OpenStudy (anonymous):

3x=11-y 3x/3=11/3-y ? @lasttccasey

OpenStudy (anonymous):

x=11/3-y ?

OpenStudy (lasttccasey):

Almost. Don't forget the y is also divided by 3. So x = 11/3 - y/3

OpenStudy (lasttccasey):

Next you will just plug this into equation two for the x variable, there by getting rid of all x's.

OpenStudy (anonymous):

-3(11/3-y/3) + 2y = 1 @lasttccasey

OpenStudy (lasttccasey):

Yes now multiply the -3 through the parenthesis, combine like terms, and solve for y. Then you're done!

OpenStudy (anonymous):

(-33/3+3/3) +2y=1 -11+2y=1 2y=1+11 2y=12 2y/2=12/2 y=6

OpenStudy (anonymous):

@lasttccasey

OpenStudy (anonymous):

But thats no answer

OpenStudy (lasttccasey):

Be careful with your math. -3(11/3-y/3) + 2y = 1 Will become: -11+ y + 2y = 1

OpenStudy (anonymous):

(-33/3+3/3) +2y=1 -11+y+2y=1 3y=12 y=4

OpenStudy (anonymous):

Thanks @lasttccasey !

OpenStudy (lasttccasey):

You're welcome! @HishamNafiseh